OpenStudy (anonymous):

(square root of)x^2+x-2

OpenStudy (anonymous):

=0

OpenStudy (anonymous):

equals 0*

OpenStudy (anonymous):

x=-2, x=1

OpenStudy (anonymous):

\[\sqrt{x^2+x-2}=0\] \[\sqrt{(x+2)(x-1)}=0\] x=-2, x=1